# UnZipLoRA: Separating Content and Style from a Single Image

### v1

@@ 3.3.2 Column Separation
Apart from learning separate concepts, we want our resulting LoRAs to be compatible, i.e. we want to be able to combine them through direct arithmetic merge to generate the subject and style together. Prompt separation effectively guides the two LoRAs to learn distinct concepts, and especially helps in achieving better destylization of the content. However, it does not guarantee compatibility since training LoRAs with different prompts can result in weight misalignment when they are combined to process the same prompt during inference. As seen in Fig. 3, output of the combined subject+style LoRA does not adhere the input prompt and overfits to the input image.
@@
Orthogonal loss. To promote compatibility between the subject and style LoRAs, we leverage the insight from ZipLoRA [32] that highly aligned LoRA columns merge poorly, and apply the following loss to $m_{c}$ and $m_{s}$:
$\vspace{-0.2em}\mathcal{L}_{\perp}=\sum_{i}|m_{c}^{i}\cdot m_{s}^{i}|,\vspace{-0.2em}$ (5)
minimizing which promotes orthogonality between the learned content and style weights. Orthogonal loss $\mathcal{L}_{\perp}$ is added to our reconstruction loss $\mathcal{L}_{DB}$ as regularizer with the weight parameter $\lambda_{\perp}$.
Figure 4: Qualitative Comparisons. We compare subject and style decomposition outputs from UnZipLoRA against DreamBooth-LoRA [29], Inspiration Tree [37], and B-LoRA [6]. Our results demonstrate that UnZipLoRA achieves superior subject and style fidelity, effectively disentangling and preserving these concepts compared to alternative approaches.
Sparse masks with importance re-calibration. To further decompose the concepts, instead of training the entire LoRA matrices, we find that training 30% of the columns in each weight matrix is sufficient to learn concepts, and it ensures weight sparsity for a good decomposition. This strategy is inspired by Liu et al. [21], who find that a small set of neurons tend to be much more salient than the others for capturing concepts.
We use a dynamic approach to select the most important 30% of the columns during training. Before the training, we initialize the column masks, $m_{s}$ and $m_{c}$, with the top 10% of the most important columns for style and content. Importance is calculated using the Cone method [21], with five warm-up training steps using the full LoRA weights. Then, during the remainder of training, for every $t$ steps, we re-calibrate the column masks by calculating the column importance of LoRA weights and adding the new top 10% of the most important columns to $m_{s}$ and $m_{c}$ until the 30% cap is reached. Our strategy brings significant improvements in compatibility of the resulting LoRAs, preventing overfitting and producing better recontextualization as shown in Fig. 3.
@@ 3.3.3 Block Separation
Furthermore, we take advantage of the analysis by the authors of B-LoRA [6], who find that certain blocks in the SDXL U-Net are more responsible for content or style. To leverage this prior knowledge, for blocks known as style-sensitive, we release the column sparsity constraint for the style LoRA,i.e., all the columns are fully trained. The same strategy is applied to the content-LoRA. This further improves the accuracy of fine details, especially for the style LoRA (see Fig. 3).
@@ 4.1 Implementation Details
Dataset. For our experiments, we use a diverse set of $40$ stylistic images representing unique styles and subjects as input images. These images are sourced from previous works, online repositories offering free-to-use content, and from state-of-the-art text-to-image generators. The attribution and license information of the input images used in our experiments is provided in the supplementary.
@@
Experimental setup. We use SDXL v1.0 [25] as our base model for all our experiments. We use rank$=64$ for our subject and style LoRAs, and fine-tune them using Adam optimizer with learning rate of $55e-$ 5. Our LoRAs are trained for $600$ steps using batch size of $1$. We keep the base SDXL model and the text encoders frozen during training. For column separation, we choose $t=200$ as the sparse column re-calibration frequency. For orthogonal loss regularization, we use $\lambda_{\perp}=0.5$ in all our experiments.
While B-LoRA inspires our block separation strategy, we diverge significantly in selecting blocks for subject and style representation. Instead of utilizing only two distinct blocks from the upsampling blocks of the U-Net, as in B-LoRA, our method leverages all upsampling blocks within the SDXL architecture for a comprehensive representation of both subject and style. We partition these upsampling blocks and assign a subset of them for content learning while the remainder is used for style learning. A detailed description of our block partitioning strategy is provided in the supplementary.
Figure 5: Recontextualization. The trained style LoRA and content LoRA can be used individually or jointly at the inference time. The learned concepts can be used to generate images in various contexts, validating our method’s robustness.
Our choices of training prompts are standard and inspired from previous works [29, 33, 32]. Our combined prompt is of the form “a <$c$> in <$s$> style”, while “A <$c$> ” and “An image in <$s$> style” are used as subject- and style-specific prompts respectively where <$c$> and <$s$> are the subject and style descriptors. As per previous works [29, 33, 32], we use subject category along with unique token (such as “sks”) to describe the subject, and use detailed artistic description for the style. For example, the prompt we use for the first input image in Fig. 4 is “A sks piano in abstract impressionism style”. More such implementation details are deferred to supplementary.
4.2 Subject and Style Decomposition
We take a single input image and apply our UnZipLoRA algorithm to obtain the style and subject LoRAs. We use these LoRAs separately to generate subject-only and style-only representations of the input image with diverse text prompts. The subject and style separation results obtained by our approach are shown in Fig. 1 and Fig. 4 from various input images. The quality of the learned concepts is further validated with various recontextualization examples in Fig 5. Our method successfully captures the intricate details of both the style and subject and can effectively disentangle both elements, as validated by the quantitative and qualitative evaluations.
@@ 4.2 Subject and Style Decomposition
Qualitative comparisons. We present qualitative comparisons of our approach with other methods in Fig. 4. We use DreamBooth-LoRA [29] as our baseline – we train subject and style LoRA models separately using the DreamBooth method with the “A <$c$> ” and “A subject in <$s$> style.” While DreamBooth-LoRA demonstrates satisfactory performance in learning the style, it fails to effectively destylize the subject and results in overfitting the input image. This highlights the inherent difficulty in disentangling subject and style from a single image, even when training LoRAs independently.
@@
UnZipLoRA is next compared with Inspiration Tree [37], a textual inversion-based concept extraction approach. Inspiration Tree aims to separate the properties of the input into a binary-tree-structured concept tree using Textual Inversion [8]. We train their model to separate subject and style with the prompt “A <$c$> in <$s$> style”, initializing each placeholder with “object” and “art”. Despite employing separate training for content and style, Inspiration Tree struggles to distinguish and disentangle these properties accurately. While it could identify the overarching category of concepts correctly, it fails to capture the intricate details of the subject or the style. This limitation stems from its reliance on Textual Inversion, which focuses on learning text embeddings rather than fine-tuning the weights, resulting in limited expressiveness and controllability.
We also compare our approach with B-LoRA [6] – a technique aimed at combining style and the subject of two different images. To adapt B-LoRA to our problem setting, we utilize the same input image to train both the subject and style representations while keeping other hyperparameters, such as the learning rate and training steps, consistent with those reported in their paper. UnZipLoRA outperforms B-LoRA in several key aspects. Firstly, UnZipLoRA demonstrates more stable content destylization and better consistency in generating variations. This is attributed to our prompt separation strategy, which ensures that each LoRA focuses exclusively on its designated concept, promoting accurate disentanglement. Secondly, UnZipLoRA exhibits superior preservation of content fidelity.
@@
As evident in the mushroom and balloon examples in Fig. 4, B-LoRA results often include residual style features and extraneous background elements, indicating overfitting and incomplete disentanglement. Furthermore, B-LoRA fails to accurately capture the color of the balloon, treating it as part of the style rather than the content. In contrast, UnZipLoRA effectively mitigates overfitting through our column separation strategy, which employs orthogonal column masking to ensure compatibility of resulting LoRA weights. The limitations of B-LoRA can be attributed to its block-wise training approach, where content information is confined to specific blocks of the U-Net, potentially leading to information loss. Moreover, unlike B-LoRA, our method trains both the LoRAs jointly, reducing the compute resource requirements significantly. More quantitative evaluation results can be found in the supplementary.
@@
Quantitative results. For quantitative comparisons, We conduct user studies comparing our method with the competing approaches. In our study, each participant is shown the input image, along with the outputs of the two methods being compared. Each output group consists of $4$ images for the subject and $4$ images for style selected randomly. Participants are asked to choose an output group that decomposes the input image into style and content more accurately. We conducted three separate user studies: one for each competing method vs. UnZipLoRA. We received $204$ responses for each study from $34$ total participants. As can be seen in Tab. 1 UnZipLoRA receives higher user preferences than all three competing methods owing to its superior disentanglement capabilities.
@@
Table 1: User Preference Study. We compare the user preference of accurate stylization and subject fidelity between our approach and competing methods. Users generally prefer our approach.
[…]
DreamBooth LoRA Inspiration Tree B-LoRA
@@
Table 2: Subject-alignment and Style-alignment Scores. We measure the cosine similarities between the input image and the decomposed output using CLIP-I and DINO features for style and subject respectively. We also provide alignment scores using CSD [34], a model more suitable for stylistic images. As indicated by higher alignment scores, UnZipLoRA provides superior subject and style fidelity on all the metrics.
DreamBooth -LoRA Inspiration Tree B-LoRA UnZipLoRA
@@
Figure 7: Ablation Examples. We show the impact of adding prompt-wise separation, column-wise separation, and block-wise separation sequentially. Each row illustrates the outputs of baseline methods and our proposed approaches, highlighting their difference in subject-style disentanglement, style fidelity, and recontextualization across different examples.
@@
We also provide quantitative comparisons of our approach using subject-alignment and style-alignment scores in Tab. 2. To calculate the alignment scores, we start with commonly used CLIP-I image embeddings for style and DINO features for objects. We calculate feature embeddings for our style and subject decomposition using CLIP-I [26] and DINO [2] features, respectively, using the input image as the reference. In both cases, we use cosine similarity as the metric and calculate averages over $8$ input images by generating $16$ samples for each. As shown in Tab. 2, UnZipLoRA achieves the highest alignment scores for both the subject and style, highlighting its superiority. DreamBooth-LoRA tends to overfit the input image, resulting in higher scores even though its performance is inferior to other methods, as depicted in qualitative results.
@@
One should note that these metrics are inherently limited, particularly in measuring style alignment, since they may not fully capture subtle stylistic nuances and can be influenced by the semantic content of the images. To this end, we use a recently proposed CSD model [34] trained specifically to extract the style descriptors from images. We use the embeddings from CSD’s content and style branches to measure subject and style fidelity respectively. The CSD alignment scores (Tab. 2) indicate a clearer separation between our method and the others, highlighting the superiority of UnZipLoRA.
Subject-style combination and recontextualization. A key advantage of UnZipLoRA is its ability to produce compatible subject and style LoRAs that can be seamlessly merged via direct addition. This allows for the generation of novel recontextualizations that faithfully incorporate both the subject and style of the original image. We demonstrate this capability in Fig. 5 through various recontextualizations, showcasing the versatility of UnZipLoRA in generating images using individual LoRAs or a combination of both. This capability to manipulate and recombine the extracted concepts underscores the robustness of our method and its potential for creative applications.
We also present additional examples of subject-style recontextualization compared with B-LoRA, which is specifically designed for it, using diverse prompts and input images in Fig. 6. The results demonstrates that our method effectively enables flexible and accurate adaptation of subjects and styles to different contexts, outperforming B-LoRA. Our generations match the requested contexts while preserving both subject and style features.
As noted in subject-only and style-only comparisons, B-LoRA suffers from overfitting and inconsistent feature preservation, issues that also affect subject-style generation. For instance, in the sun hat example of Fig. 6, B-LoRA fails to retain key features of the input hat, demonstrating these limitations. Moreover, our method demonstrates clear superiority over B-LoRA in handling diverse scenarios, including those involving multiple subjects or complex relationships. For instance, in the pizza example in Fig. 6, B-LoRA incorrectly applies subject features to the cat, generating pizza-like textures. In contrast, our method perfectly preserves the cat and the pizza, generating a contextually appropriate output.
4.3 Ablation Examples
In this section, we present qualitative examples to illustrate the impact of the different components of our method. We use DreamBooth [29] as the baseline, and iteratively add prompt-, column-, and block-separation strategies. We compare the performance of each ablated version of our method to demonstrate effectiveness of each component. For simplicity, let us refer to ablations of our method as M1, M2, and M3, where:
  • M1 is Baseline + Prompt Separation
  • M2 is Baseline + Prompt + Column Separation
  • M3 is our full UnZipLoRA method (Baseline + Prompt + Column + Block Separation)
We provide a qualitative comparison of the effects of each proposed component in Sec. 3.3 in Fig. 7. Each row demonstrates the recontextualization of subject, style and combined subject-style across different configurations.
As shown in Fig. 7, adding only the prompt-wise separation (M1) to the baseline significantly improves the disentanglement of content with style compared to baseline. However, M1 still struggles to capture the style complexities from the input images, and shows strong overfitting to the input image in subject-style recontextualization highlighting the incompatibility of resulting LoRAs.
Next, as described in detail in Sec. 3, incorporating column-wise separation (M2) reduces the interference between subject and style LoRAs and improves their compatibility while preserving subject and style generation abilities. For example, details like the color of the microscope in Fig. 7 are retained, while the combined recontextualization performance improves significantly. However, improvements in style decomposition remain modest, indicating that using a small portion of the columns for style is insufficient for effective style learning.
Furthermore, we introduce block separation (M3) with subject- and style-specific blocks, further refining the disentanglement of subject and style. This addition significantly enhances the model’s ability to capture fine-grained stylistic features owing to dedicated style-learning blocks. Fig. 7 shows that all three separation strategies together achieves the best results across all examples, enabling comprehensive disentanglement of subject and style and flexible recontextualization.
4.4 Limitations
In some cases involving abstract artistic styles, UnZipLoRA may not achieve perfect destylization. For instance, in Fig. 8, the kangaroo example exhibits artifacts on its leg due to the influence of the curved line-art style, while the bear’s form appears slightly unnatural due to the high level of abstraction in the original style. This highlights the challenge of disentangling content from highly abstract styles where the style significantly influences the subject’s depiction.

### v2

@@ 3.3.2 Column Separation
Apart from learning separate concepts, we want our resulting LoRAs to be compatible, i.e. we want to be able to combine them through direct arithmetic merge to generate the subject and style together. Prompt separation effectively guides the two LoRAs to learn distinct concepts, and especially helps in achieving better destylization of the content. However, it does not guarantee compatibility since training LoRAs with different prompts can result in weight misalignment when they are combined to process the same prompt during inference.
@@
Figure 3: Recontextualization. The trained style and content LoRAs can be used individually or jointly at the inference time. The learned concepts can be used to generate images in various contexts, validating our method’s robustness. Additional examples in the supplementary.
Figure 4: Qualitative Comparisons. Subject and style decomposition outputs from UnZipLoRA compared against DreamBooth-LoRA [30], Inspiration Tree [38], and B-LoRA [6]. UnZipLoRA achieves superior subject and style fidelity, disentangling and preserving these concepts more effectively. More examples can be found in the supplementary.
Sparse masks with importance re-calibration. To further decompose the concepts, instead of training the entire LoRA matrices, we find that training only a fraction of total columns in each weight matrix is sufficient to learn the concepts, and it ensures weight sparsity for improved decomposition. This strategy is inspired by Liu et al. [22], who find that a small set of neurons tend to be much more salient than the others for capturing concepts.
We use a dynamic approach to select the most important $N\%$ of the columns during training. Before the training, we initialize the column masks, $m_{s}$ and $m_{c}$, with the top $(N/3)\%$ of the most important columns for style and content. Importance is calculated using the Cone method [22], with five warm-up training steps using the full LoRA weights. Then, during the remainder of training, for every $t$ steps, we re-calibrate the column masks by calculating the column importance of LoRA weights and adding the new top $(N/3)\%$ of the most important columns to $m_{s}$ and $m_{c}$ until the $N\%$ cap is reached. In practice, $N$ in the range of $25$ to $40$ work well, and we choose $N=30$ for our experiments.
Orthogonal loss. To promote compatibility between the subject and style LoRAs, we leverage the following orthogonality loss on $m_{c}$ and $m_{s}$:
$\vskip-1.99997pt\mathcal{L}_{\perp}=\sum_{i}|m_{c}^{i}\cdot m_{s}^{i}|,\vskip-1.99997pt$ (5)
minimizing which promotes orthogonality between the learned content and style weights [33]. Orthogonal loss $\mathcal{L}_{\perp}$ is added to our reconstruction loss $\mathcal{L}_{DB}$ as regularizer with the weight parameter $\lambda_{\perp}$.
This strategy brings significant improvements in compatibility of the resulting LoRAs, preventing overfitting to the input image and producing better recontextualizations, as shown in Fig. 5 (row M2).
@@ 3.3.3 Block Separation
B-LoRA [6] showed that certain blocks in the SDXL U-Net are more responsible for content and some are more responsible for style. We can leverage this insight by relaxing the column sparsity constraints on the style and subject LoRAs corresponding to the style-sensitive and subject-sensitive blocks, respectively. In other words, all the LoRA columns in these blocks are fully trained without sparse masks. In our attempt to take advantage of B-LoRA’s block constraints, we find their split between subject and style is coarse-grained: tellingly, they fail to extract and represent the subject from the input image in its unstylized appearance. Therefore, we extend their approach by involving more blocks, and performing more fine-grained block-wise allocations of subject and style within the Up-blocks of SDXL U-Net (details in supplementary). This further improves the accuracy of fine details, especially for the style LoRA (see last row of Fig. 5).
@@ 4.1 Implementation Details
Dataset. In our experiments, we use a set of $40$ diverse images with unique styles and subjects. These are collected from previous work, state-of-the-art text-to-image generators, and open-source repositories. Attribution information is provided in the supplementary material.
@@
Experimental setup. We use SDXL v1.0 [26] as our base model. Subject and style LoRAs are trained with rank =64=64 using Adam (learning rate = 5e55e-5) for $600$ steps with batch size =1=1, keeping the base SDXL weights and text encoders frozen. Column separation uses $t=200$, $N=30\%$, and weight of orthogonal loss is set to $\lambda_{\perp}=0.5$ in all our experiments. Our block separation uses all upsampling blocks in the SDXL U-Net, unlike B-LoRA that uses just two. See supplementary material for details of which blocks we assign to content and which ones to style learning.
4.2 Qualitative Results
As shown in Fig. 1, the subject and style LoRAs obtained by UnZipLoRA can be used separately to generate new representations of subject-only and style-only concepts. A key advantage of UnZipLoRA is its ability to produce compatible subject and style LoRAs that can be seamlessly merged via direct addition. This allows for the generation of novel recontextualizations that faithfully incorporate both the subject and style of the original image. Fig. 3 demonstrates this capability through recontextualizations using either individual LoRAs or a combination of both. More results and comparisons are included in the supplementary material.
4.3 Comparative Evaluation
In this section, we compare UnZipLoRA with three recent methods: DreamBooth-LoRA [30], Inspiration Tree [38], and B-LoRA [6]. As shown in Fig. 4, our results are clearly superior qualitatively, and as a consequence, they are strongly preferred in a user study (Tab. 1).
@@ 4.3 Comparative Evaluation
For DreamBooth-LoRA [30], we train subject and style LoRA models separately using the DreamBooth method with the trigger phrases “A <cc>” and “A <cc> in <ss> style.” Note that the <cc> and <ss> used here are identical to those in UnZipLoRA. As can be seen from Fig. 4, although DreamBooth-LoRA demonstrates satisfactory performance in learning the style, it fails to effectively destylize the subject and results in overfitting. This highlights the inherent difficulty in disentangling subject and style from a single image, even when training the LoRAs independently.
@@
Inspiration Tree [38] is a concept extraction method that decomposes the input into a binary concept tree using Textual Inversion [8]. We train their model to extract subject and style using the prompt “A <cc> in <ss> style”, and follow their default setup that initializes the placeholders <cc> and <ss> with “object” and “art” respectively. Despite employing separate training for content and style, Inspiration Tree struggles to disentangle them accurately. While it can identify the overarching category of concepts correctly, it fails to capture the intricate details of the subject or the style. This limitation stems from its reliance on Textual Inversion, which focuses on learning text embeddings rather than fine-tuning the weights, resulting in limited expressiveness and controllability.
@@
We also compare our approach with B-LoRA [6] – a technique aimed at combining style and subject of two different images. We follow the default setup of B-LoRA, and independently train both the subject and style representations on the same input image using same trigger phrases as UnZipLoRA. We keep other hyperparameters (learning rate, training steps, etc.) consistent with  [6]. As evident in the lantern and balloon examples in Fig. 4, B-LoRA outputs often retain residual style features and extraneous background elements, indicating overfitting and incomplete disentanglement. Furthermore, B-LoRA fails to accurately capture the color of the balloon, treating it as part of the style rather than the content. The limitations of B-LoRA can be attributed to its block-wise training approach, where content information is confined to specific blocks of the U-Net, potentially leading to information loss. By contrast, UnZipLoRA demonstrates more stable content destylization and better consistency in generating variations. Moreover, unlike B-LoRA, our method trains both the LoRAs jointly, reducing the compute requirements significantly.
@@
Tab. 1 presents results of user studies comparing our method with the competing approaches for both subject/style decomposition and recontextualization. In our study, each participant is shown the input image, along with the outputs of two methods being compared (the methods are not labeled and their order is arbitrary). Each output group consists of $4$ images for the subject and $4$ images for style selected randomly, and participants are asked to choose an output group that decomposes the input image into style and content more accurately (see supplementary material for an example screenshot of the interface). We conducted three separate user studies – one for each competing method vs. UnZipLoRA – and received $204$ responses for each from $34$ total participants. As can be seen in Tab. 1, UnZipLoRA is strongly preferred over all three competing methods for both decomposition and recontextualization.
@@
Table 1: User Preference Study. We compare the user preference for subject-style decomposition (top), and combined subject-style recontextualization (bottom) between our approach and competing methods. Users generally prefer our approach in both. We received $204$ responses for each study from $34$ total participants.
$\%$ Preference for UnZipLoRA over:
DreamBooth LoRA
Inspiration Tree
B-LoRA
Decomposition
@@
Recontextualization
$98.10\%$ $9.17\% 77$$7.14\%$
Table 2: Subject-alignment and Style-alignment Scores. Comparisons for Content and Style Decomposition.
DB-LoRA
Insp. Tree
B-LoRA
UnZipLoRA
@@
Tab. 2 further presents comparisons using automatic subject-alignment and style-alignment scores. We calculate alignment scores between decomposition output and the input image using standard CLIP-I image embeddings [27] for style and DINO features [2] for content. In both cases, we use cosine similarity as the metric and calculate averages over $8$ input images by generating $16$ samples for each. As can be seen from Tab. 2, UnZipLoRA achieves the highest alignment scores for both subject and style, highlighting its superiority. DreamBooth-LoRA tends to overfit the input image, resulting in higher scores even though its outputs are qualitatively inferior to those of other methods.
@@
CLIP-I and DINO metrics are inherently limited, especially in measuring style alignment, since they may not fully capture stylistic nuances and can be influenced by the semantic content of the images. As a more sensitive metric, we use a recently proposed CSD model [35] trained specifically to extract the style descriptors from images. We use the embeddings from CSD’s content and style branches to measure subject and style fidelity respectively. The CSD alignment scores (third and fourth line of Tab. 2) yield a clearer separation between our method and the others.
4.4 Ablation Study
In this section, we present an ablation study to justify the effectiveness of our system design. As illustrated in Fig. 5, we start from DreamBooth [30] as the baseline and successively add the key components of our method: prompt separation (M1), column separation (M2), and block separation (M3). Tab. 3 reports user preferences for each model version over the previous one (based on $204$ responses from $34$ participants).
As can be seen from Fig. 5, the DreamBooth baseline is incapable of subject-style disenganglement and has very weak recontextualization ability. Adding prompt separation (M1) to the baseline enables successful extraction of a realistic content LoRA and improves recontextualization. However, prompt separation alone struggles to capture all the complexities of the style from the input image (only 12.35% user preference over baseline for style decomposition), and shows strong overfitting to the input image in combined recontextualization, indicating incompatibility of the learned LoRAs. Adding column separation (M2) on top of M1 reduces the interference between subject and style LoRAs while also improving style generation abilities. For example, details like the color of the microscope in Fig. 5 are retained, and the combined recontextualization performance improves significantly. However, improvements in style decomposition remain modest, indicating that using a small portion of the columns for style is insufficient for effective style learning. Finally, adding block separation (M3) with subject- and style-specific blocks on top of M2 significantly enhances the model’s ability to capture fine-grained stylistic features. The last row of Fig. 5 shows that using all three separation strategies together achieves the best results, enabling comprehensive disentanglement of subject and style and flexible recontextualization.
Table 3: Ablation User Study. Baseline: Dreambooth-LoRA; M1: Prompt-separation only; M2: Prompt- and column- separations; M3 (full): Prompt-, column-, and block- separations.
$\%$ Preference for:
M1 over Baseline M2 over M1 M3 over M2
[…]
Style Decomposition $12.35\%$ $39.51\%$ $\mathbf{86.42\%}$
Combined Recontextualization $92.80\%$ $93.64\%$ $\mathbf{61.90\%}$
Figure 5: Illustration of ablations. Each column shows the performance of subject LoRA, style LoRA, and combined LoRA.
4.5 Applications and extensions
Cross-combination of subject and style LoRAs from different images. The subject and style LoRAs produced by UnZipLoRA open up a possibility for cross-combination: pairing a subject LoRA from one image with a style LoRA from another. Fig. 6 shows such cross-combination results where the LoRAs are combined by direct addition. While these LoRAs are not explicitly trained together (and thus not subject to the orthogonality constraints enforced by ZipLoRA [33]), the inherent separation imposed by our column and block strategies generally results in higher compatibility than generic DreamBooth-LoRAs trained without such constraints. Consequently, direct arithmetic merger yields promising cross-stylization results.
Figure 6: UnZipLoRA provides the valuable capability of cross-composition using subject and style LoRAs from different images.
Extension to other architectures. To demonstrate the possibility of extending our method beyond SDXL, we train on KOALA [19], a more efficient recent model with a leaner U-Net architecture. As shown in Fig. 7, our method, when applied to KOALA, accurately captures subject and style and allows for successful recontextualization (though the overall quality of the results is not as high as for SDXL due to limited capacity and lower parameter count of KOALA). The core idea of UnZipLoRA, that of simultaneously training two LoRAs on the same input image, should also be applicable to other architectures like DiT [25], though this extension is beyond the scope of present work.
Figure 7: Our approach generalizes effectively, as demonstrated by successful results on the more recent KOALA diffusion model.